OpenStudy (anonymous):

Can you please help me solve -2/3(x-4)>5/6(x+1)

OpenStudy (chriss):

I plugged it into Maple, I got\[x <\frac{11}{9}\] If you plug that in for x it works so the solution is correct. I'm having a little trouble working through the algebra on it, but if you need the steps I'll keep plugging away at it til I figure it out.

OpenStudy (anonymous):

Chriss OP probably meant this : $$\frac{-2}{3 (x - 4)} > \frac{5}{6 (x + 1)} $$

OpenStudy (anonymous):

which implies $$ x<-1\left\|\frac{16}{9}<x<4\right.$$

OpenStudy (chriss):

\[-\frac{2}{3}x+\frac{8}{3}>\frac{5}{6}x+\frac{5}{6}\] \[-\frac{2}{3}x+\frac{16}{6}>\frac{5}{6}x+\frac{5}{6}\] \[-\frac{2}{3}x+\frac{11}{6}>\frac{5}{6}x\] \[\frac{11}{6}>\frac{9}{6}x\] \[\frac{11}{6}*\frac{6}{9}=x\] \[\frac{66}{54}=\frac{11}{9}=x\]

OpenStudy (chriss):

@Fool that may be. I see how it could be read either way. In the event that my interpretation of the way it is written is correct though, there is the solution.